# P2 Law: Scaling Law for Post-Training After Model Pruning

### v2

@@ Abstract
Xiaodong Chen1,2, Yuxuan Hu1,2, Xiaokang Zhang1,2, Yanling Wang3   Cuiping Li1,2Hong Chen1,2Jing Zhang1,2thanks: Corresponding author. 1School of Information, Renmin University of China, Beijing, China 2Key Laboratory of Data Engineering and Knowledge Engineering, Beijing, China 3 Zhongguancun Laboratory, China {chenxiaodong,huyuxuan1999,zhang-jing,zhang2718,licuiping,chong}@ruc.edu.cn  wangyl@zgclab.edu.cn
@@ 1 Introduction
Figure 1: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series models pruned by depth pruning with a pruning rate of approximately 15%. $C$ denotes the computational cost, which is calculated by $C=6ND$ Kaplan et al. (2020), where $N$ denotes the model size after pruning, and $D$ denotes the number of post-training tokens.
@@
To address the problem, we conduct pilot experiments on the Llama-3 Dubey et al. (2024) and Qwen-2.5 series models Team (2024), applying both typical structured and semi-structured pruning methods. In specific, we observe several trends in the post-training loss curves, allowing us to identify the necessary conditions that the scaling Law for Post-training after model Pruning (P2 Law) must satisfy. Building on the Chinchilla scaling law Hoffmann et al. (2022) proposed for pre-training and the identified conditions, we define multiple parameterizations of our P2 Law and select the most suitable parameterization. To assess the fit of different parameterizations to P2 Law, we introduce a new metric named Average Slope Difference (ADS). As scaling laws are used to find the suitable training data size by balancing cost and performance, focusing on the slope of the predicted loss curve rather than the predicted loss values, the ADS metric is designed to measure the slope discrepancy between predicted and actual loss curves. Finally, P2 Law is parameterized as,
@@
In this paper, we conduct a series of experiments to validate the P2 Law. Taking Llama-3 series models pruned by depth pruning with a pruning rate of approximately 15% as an example, Figure 1 illustrates that P2 Law accurately predicts the actual post-training losses of the pruned model checkpoints, where $C$ represents the computational cost calculated as $C=6ND$ Kaplan et al. (2020), $N$ is the model size after pruning, and $D$ is the number of post-training tokens. Utilizing the post-training loss curves derived by P2 Law, we can accurately predict that the computational cost required for the post-training loss of Llama-3.2-1B to start decreasing gently is approximately $10^{4}$. This predicted size of post-training data provides a good balance between cost and performance. Furthermore, we evaluate the generalization ability of the P2 Law, demonstrating that P2 Law can effectively generalizes to larger dataset sizes, larger model sizes, and higher pruning rates.
@@
  • We propose ADS, an effective metric for the evaluation of parameterizations of scaling laws for the post-training of pruned LLMs.
@@ 2.2 Post-Training
After the pruning, we conduct post-training on the pruned LLMs to mitigate the performance decline. For LLMs pruned using depth or width pruning, we train all parameters of the pruned LLMs. For sparse LLMs derived from 2:4 semi-structured pruning, inspired by SP-LoRA Anonymous (2024), we combine the updated weight from each training iterate with the sparse mask during the post-training process to ensure the model’s sparsity, further post-training details about the 2:4 semi-structured pruning is provided in Appendix B.3.
@@ 3.1 Settings
Post-Training. For Llama-3.2-3B, Qwen-2.5-3B and Llama-3.1-8B, we randomly select 1B tokens from SlimPajama for post-training. For Llama-3.2-1B, Qwen-2.5-0.5B, Qwen-2.5-1.5B, we randomly select 0.5B tokens from SlimPajama for post-training. During the post-training process, we set the learning rate to 2e-5 and the batch size to 256k tokens. All post-training processes are conducted on 4 Nvidia A800-80G GPUs and 4 Nvidia A6000-48G GPUs. The entire training process takes a total of 500 hours.
@@ 3.2 Trends of the Post-Training Loss Curves
In Figures 2 and 4, we present the post-training loss curves for the Llama-3 series models pruned by depth pruning and 2:4 semi-structured pruning. Additional post-training loss curves (exhibiting similar trends) are shown in Figures 101112, and 14 in the Appendix C. By analyzing the post-training loss curves, we observe the following trends:
@@ 3.3 Necessary Conditions Satisfied by P2 Law
  • Condition 2. As derived from Trend 1 in Section 3.2, under similar pruning rates, the post-training loss curves of smaller LLMs decrease faster as the number of post-training tokens increases:
@@
  • Condition 3. From Eq.4 in Trend 2, the relative post-training loss $\Delta\mathcal{L}$ follows a power-law relationship with the pruning rate $\rho$.:
@@ 4.1 Metric for Accessing Law Fitting
Following prior work Que et al. (2024), we utilize both $R^{2}$ Fisher (1922) and Huber loss Huber (1992) to evaluate different parameterizations of scaling law. The $R^{2}$ value, or coefficient of determination, measures the fitting quality, with values closer to 1 indicating a better fit. Huber loss, a robust loss function, blends the characteristics of mean squared error and mean absolute error, making it less sensitive to outliers. The Huber loss is a positive number, and a lower Huber loss suggests a better fit.
Scaling laws are often used to determine the optimal amount of training data by balancing computational cost and model performance. For instance, as shown in Figure 5, there is one actual loss curve and two predicted loss curves. Traditional metrics like $R^{2}$ and Huber loss indicate that predicted curve 1 better matches the actual curve. However, when predicted curve 1 is used to predict the point where the actual curve starts to flatten, it results in a large error. On the other hand, while predicted curve 2 deviates more from the actual curve in terms of absolute values, its slope is consistently closer to the actual curve. This makes its prediction of the flattening point much more accurate. To address this issue, we propose a new metric called Average Slope Difference (ASD), which measures the difference between the slope of the loss curve predicted by the scaling law and the slope of the actual loss curve. ASD is formally defined as:
@@
Figure 5: A example showcasing the advantages of ADS. The ADS of predicted loss curve 2 is lower because its slope is closer to that of the actual loss curve.
@@
Figure 6: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series model pruned by depth pruning method.
@@ 4.2 Derivation of P2 Law
where $N_{C}$, $D_{C}$, $E$, $\alpha$, and $\beta$ are constants, $N$ represents the model size, $D$ denotes the number of pre-training tokens and $\mathcal{L}$ represents the model’s loss. Compared to the OpenAI scaling law Kaplan et al. (2020), the Chinchilla scaling law demonstrates superior performance (detailed in Appendix D). Therefore, we adopt the Chinchilla scaling law as the foundational parameterization for our P2 Law. Combining the pruning rate $\rho$ and the model’s loss $\mathcal{L}_{0}$ before pruning, we define the following three candidate parameterizations:
$\begin{split}\mathcal{L}_{1}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (10)
$\begin{split}\mathcal{L}_{2}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (11)
$\begin{split}\mathcal{L}_{3}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{D_{C}}{D^{\beta}}+E)\end{split}$ (12)
where $N_{C}$, $D_{C}$, $E$, $\alpha$, $\beta$, $\gamma$ and $\delta$ are constants, $N_{0}$ denotes the model size before pruning, $D$ denotes the number of post-training tokens and $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$ denote the pruned model’s post-training loss. Additionally, since there is no pruning rate in the 2:4 semi-structured pruning, P2 Law for the 2:4 semi-structured pruning does not need to satisfy Condition 3, and we adjust the parameterizations to:
@@
Figure 7: Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-3B pruned by width pruning).
Figure 8: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Qwen-2.5 series models pruned by width pruning).
Figure 9: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models pruned by width pruning).
We utilize all checkpoints to fit the three candidate parameterizations through Levenberg-Marquardt method Moré (2006), and the specific parameter values (i.e., the values of $N_{c}$, $D_{c}$, $E$, $\alpha$, $\beta$, $\gamma$, and $\delta$) for the fitted $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ are provided in Table 5 in Appendix E. As shown in Table 2, $\mathcal{L}_{1}$ significantly outperforms $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ in terms of the $R^{2}$, Huber loss, and ASD metrics. Additionally, as shown in Table 6 in Appendix E, after our calculation and verification, $\mathcal{L}_{2}$ and some fitted $\mathcal{L}_{3}$ fails to satisfy Condition 2. In contrast, all of the fitted $\mathcal{L}_{1}$ satisfy all three conditions. Based on the experimental results, we select $\mathcal{L}_{1}$ as the parameterization for our P2 Law.
In Figure 6, we show the loss curves $\mathcal{L}_{1}$ derived by P2 Law alongside the actual checkpoints of the Llama-3 series models pruned by depth pruning, where the computational cost ($C$) is approximated using the empirical formula $C=6ND$ Kaplan et al. (2020), and $N$ denotes the model size after pruning. Additional loss curve derived by P2 Law are shown in Figure 151617 and 18 in Appendix F. As shown in these figures, the loss curve derived by P2 Law accurately aligns with all actual checkpoints, under all the three pruning methods, except for the width pruning on Llama-3.1-8B. As shown in Figure 2(c) and 10(c), for Llama-3.1-8B, we observe that depth pruning outperforms width pruning at similar pruning rates, which contrasts with the observations in other cases. This suggests that width pruning on Llama-3.1-8B may lead to anomalous performance, making our law unsuitable for this special scenario. We elaborate on this anomalous performance of width pruning on Llama-3.1-8B further in Appendix G.
@@ 4.3.1 Settings
Model size. We fit the P2 Law using checkpoints from smaller LLMs, including Llama-3.2-1B, Llama-3.2-3B, Qwen-2.5-0.5B, and Qwen-2.5-1.5B. We then validate the fitted P2 Law on checkpoints from larger LLMs, including Llama-3.1-8B and Qwen-2.5-3B. Taking the Qwen-2.5 series models as an example, we fit the P2 Law using all checkpoints from Qwen-2.5-0.5B and Qwen-2.5-1.5B, and subsequently validate it with the actual checkpoints of Qwen-2.5-3B across three pruning rates. Due to the limited number of available actual loss curves for 2:4 semi-structured pruning, we did not conduct experiments for this method.
@@
Pruning Rate. We fit the P2 Law using checkpoints from lower pruning rates and then validate it with checkpoints from higher pruning rate. Taking width pruning of the Qwen-2.5 series models as an example, we fit the P2 Law using checkpoints from these models at lower pruning rates (0.15 and 0.25) and then validate it with the actual checkpoints at a higher pruning rate of 0.35. Since there is no pruning rate in the 2:4 semi-structured pruning, we only explore the generalization ability on pruning rates under depth pruning and width pruning.
@@
Due to the anomaly of the width pruning on Llama-3.1-8B (detailed in the last paragraph of Section 4.2), we do not consider the generalization experiments on this model with width pruning.
4.3.2 Dataset Size Generalization
@@
The evaluation results are shown in Table 3, and the loss curves of Qwen-2.5-3B (pruned by width pruning) derived by P2 Law are illustrated in Figure 9. Additional loss curves derived by P2 Law are provided in Figures 202122, and 23 in Appendix H. The results in Table 3 show that the loss curves derived by P2 Law accurately matches the validation checkpoints, indicating that the P2 Law generalizes well to larger dataset sizes.
4.3.3 Model Size Generalization
@@
The evaluation results are presented in Table 3, and the loss curves of Qwen-2.5-3B predicted by P2 Law (pruned by width pruning) are visualized in Figure 9. Additional loss curves predicted by P2 Law are shown in Figure 25 in Appendix H. As shown in Table 3, the P2 Law fitted on smaller LLMs performs poorly in $R^{2}$ and Huber loss when applied to larger models, indicating challenges in generalizing to larger, unseen models. However, the low ASD suggests it still captures the slope of the actual loss curve. This trend is also seen in Figure 25, where despite a gap between predicted and actual loss curves, the predicted and actual loss curves align in their downward trend after training stabilizes. This suggests P2 Law fitted from smaller LLMs can still predict the optimal computation cost point for larger LLMs, confirming its generalization feasibility.
4.3.4 Pruning Rate Generalization
@@
We present the generalization evaluations in Table 3 and illustrate the loss curves of Qwen-2.5 series models predicted by P2 Law (pruned by width pruning) in Figure 9. Additional loss curves predicted by P2 Law are provided in Figure 26 and 27 in Appendix H. As shown in the Figure 9 and Table 3, the values of different metrics indicate that the actual loss curves closely align with the predicted loss curves, suggesting that the P2 Law generalizes well to higher pruning rates.

### v3

@@ Abstract
Xiaodong Chen2,3, Yuxuan Hu2,3thanks: Xiaodong Chen and Yuxuan Hu have equal contribution., Xiaokang Zhang2,3, Yanling Wang4   Cuiping Li1,2Hong Chen1,2Jing Zhang1,2thanks: Corresponding author. 1Engineering Research Center of Database and Business Intelligence, MOE, China 2School of Information, Renmin University of China,Beijing, China 3Key Laboratory of Data Engineering and Knowledge Engineering, MOE, China 4 Zhipu AI, China {chenxiaodong,huyuxuan1999,zhang-jing,zhang2718,licuiping,chong}@ruc.edu.cn  wangyl@zgclab.edu.cn
@@ 1 Introduction
Figure 1: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series models pruned by depth pruning with a pruning rate of approximately 15%. Compute ($C$) denotes the computational cost, which is calculated by $C=6ND$ Kaplan et al. (2020), where $N$ denotes the model size after pruning, and $D$ denotes the number of post-training tokens.
@@
To address the problem, we conduct pilot experiments on the Llama-3 Dubey et al. (2024) and Qwen-2.5 series models Team (2024), applying both typical structured and semi-structured pruning methods. In specific, we observe several trends in the post-training loss curves, allowing us to identify the necessary conditions that the scaling Law for Post-training after model Pruning (P2 Law) must satisfy. Building on the Chinchilla scaling law Hoffmann et al. (2022) proposed for pre-training and the identified conditions, we define multiple parameterizations of our P2 Law and select the most suitable parameterization. To assess the fit of different parameterizations to P2 Law, we introduce a new metric named Average Slope Difference (ASD). As scaling laws are used to find the suitable training data size by balancing cost and performance, focusing on the slope of the predicted loss curve rather than the predicted loss values, the ASD metric is designed to measure the slope discrepancy between predicted and actual loss curves. Finally, P2 Law is parameterized as,
@@
In this paper, we conduct a series of experiments to validate the P2 Law. Taking Llama-3 series models pruned by depth pruning with a pruning rate of approximately 15% as an example, Figure 1 illustrates that P2 Law accurately fits the actual post-training losses of the pruned model checkpoints, where compute ($C$) represents the computational cost calculated as $C=6ND$ Kaplan et al. (2020), $N$ is the model size after pruning, and $D$ is the number of post-training tokens. Utilizing the post-training loss curves derived by P2 Law, we can accurately predict that the computational cost required for the post-training loss of Llama-3.2-1B to start decreasing gently is approximately $10^{4}$. This predicted size of post-training data provides a good balance between cost and performance. Furthermore, we evaluate the generalization ability of the P2 Law, demonstrating that P2 Law can effectively generalizes to larger dataset sizes, larger model sizes, and higher pruning rates.
@@
  • We propose ASD, an effective metric for the evaluation of parameterizations of scaling laws for the post-training of pruned LLMs.
@@ 2.2 Post-Training
After the pruning, we conduct post-training on the pruned LLMs to mitigate the performance decline. For LLMs pruned using depth or width pruning, we train all parameters of the pruned LLMs. For sparse LLMs derived from 2:4 semi-structured pruning, inspired by LoRS Hu et al. (2025), we combine the updated weight from each training iterate with the sparse mask during the post-training process to ensure the model’s sparsity, further post-training details about the 2:4 semi-structured pruning is provided in Appendix B.3.
@@ 3.1 Settings
Post-Training. For Llama-3.2-3B, Qwen-2.5-3B and Llama-3.1-8B, we randomly select 1B tokens from SlimPajama for post-training. For Llama-3.2-1B, Qwen-2.5-0.5B, Qwen-2.5-1.5B, we randomly select 0.5B tokens from SlimPajama for post-training. During the post-training process, we set the learning rate to 2e-5 and the batch size to 262k tokens. All post-training processes are conducted on 4 Nvidia A800-80G GPUs and 4 Nvidia A6000-48G GPUs. The entire training process takes a total of 500 hours. For more details about batch size and learning rate settings, please refer to the Appendix C.
@@ 3.2 Trends of the Post-Training Loss Curves
In Figures 2 and 4, we present the post-training loss curves for the Llama-3 series models pruned by depth pruning and 2:4 semi-structured pruning. Additional post-training loss curves (exhibiting similar trends) are shown in Figures 8910, and 12 in the Appendix D. By analyzing the post-training loss curves, we observe the following trends:
@@ 3.3 Necessary Conditions Satisfied by P2 Law
  • Condition 2. As derived from Trend 1 in Section 3.2, under similar pruning rates, the post-training loss curves of smaller LLMs decrease faster as the number of post-training tokens $D$ increases:
@@
  • Condition 3. From Eq.4 in Trend 2, the relative post-training loss $\Delta\mathcal{L}$ follows a power-law relationship with the pruning rate $\rho$:
@@ 4.1 Metric for Accessing Law Fitting
Following prior work Que et al. (2024), we utilize both $R^{2}$ Fisher (1922) and Huber loss Huber (1992) to evaluate different parameterizations of scaling law. The $R^{2}$ value, reflecting the proportion of variance explained, trends toward 1 as the fit becomes more robust. Huber loss, a robust loss function, blends the characteristics of mean squared error and mean absolute error, making it less sensitive to outliers. The Huber loss is a positive number, and a lower Huber loss suggests a better fit.
Scaling laws are often used to determine the optimal amount of training data by balancing computational cost and model performance. For instance, as shown in Figure 5, there is one actual loss curve and two predicted loss curves. Traditional metrics like $R^{2}$ and Huber loss indicate that predicted curve 1 better matches the actual curve. However, the convergence trend predicted by curve 1 deviates significantly from the actual convergence trend. While curve 1 predicts that the loss has flattened, the actual loss continues to decrease. On the other hand, while predicted curve 2 deviates more from the actual curve in terms of absolute values, its slope is consistently closer to the actual curve. This makes its prediction of the flattening point much more accurate. To address this issue, we propose a new metric called Average Slope Difference (ASD), which measures the difference between the slope of the loss curve predicted by the scaling law and the slope of the actual loss curve. ASD is formally defined as:
@@
Figure 5: An example showcasing the advantages of ASD. The ASD of predicted loss curve 2 is lower because its slope is closer to that of the actual loss curve.
@@
Figure 6: Loss curves derived by P2 Law and the actual checkpoints of Llama-3 series models pruned by depth pruning.
@@ 4.2 Derivation of P2 Law
where $N_{C}$, $D_{C}$, $E$, $\alpha$, and $\beta$ are constants, $N$ represents the model size, $D$ denotes the number of pre-training tokens and $\mathcal{L}$ represents the model’s loss. Compared to the OpenAI scaling law Kaplan et al. (2020), the Chinchilla scaling law demonstrates superior performance (detailed in Appendix E). Therefore, we adopt the Chinchilla scaling law as the foundational parameterization for our P2 Law. Combining the pruning rate $\rho$ and the model’s loss $\mathcal{L}_{0}$ before pruning, we define the following three candidate parameterizations:
$\begin{split}\mathcal{L}_{1}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$
$\begin{split}\mathcal{L}_{2}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{N_{C}}{N_{0}^{\alpha}}+\frac{D_{C}}{D^{\beta}}+E)\end{split}$
$\begin{split}\mathcal{L}_{3}(N_{0},D,\rho,\mathcal{L}_{0})=\mathcal{L}_{0}+(\frac{1}{\rho})^{\gamma}(\frac{1}{N_{0}})^{\delta}(\frac{D_{C}}{D^{\beta}}+E)\end{split}$
where $N_{C}$, $D_{C}$, $E$, $\alpha$, $\beta$, $\gamma$ and $\delta$ are constants, $N_{0}$ denotes the model size before pruning, $D$ denotes the number of post-training tokens and $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$ denote the pruned model’s post-training loss. Additionally, since there is no pruning rate in the 2:4 semi-structured pruning, P2 Law for the 2:4 semi-structured pruning does not need to satisfy Condition 3. As a result, both the pruning rate and the loss before pruning are omitted and we adjust the parameterizations to:
@@
(a) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation.
(b) P2 Law is fitted using checkpoints from smaller LLMs and used to predict the loss curves of larger LLMs.
(c) P2 Law is fitted using checkpoints from smaller pruning rates and used to predict the loss curves of larger ones.
Figure 7: Generalization of the P2 Law for Qwen-2.5 series models pruned by width pruning.
We utilize all checkpoints to fit the three candidate parameterizations through Levenberg-Marquardt method Moré (2006), and the specific parameter values (i.e., the values of $N_{C}$, $D_{C}$, $E$, $\alpha$, $\beta$, $\gamma$, and $\delta$) for the fitted $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ are provided in Table 5 in Appendix F. As shown in Table 2, $\mathcal{L}_{1}$ significantly outperforms $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ in terms of the $R^{2}$, Huber loss, and ASD metrics. Additionally, as shown in Table 6 in Appendix F, after our calculation and verification, $\mathcal{L}_{2}$ and some fitted $\mathcal{L}_{3}$ fails to satisfy Condition 2. In contrast, all of the fitted $\mathcal{L}_{1}$ satisfy all three conditions. Based on the experimental results, we select $\mathcal{L}_{1}$ as the parameterization for our P2 Law.
In Figure 6, we show the loss curves $\mathcal{L}_{1}$ derived by P2 Law alongside the actual checkpoints of the Llama-3 series models pruned by depth pruning, where the compute ($C$) is approximated using the empirical formula $C=6ND$ Kaplan et al. (2020), and $N$ denotes the model size after pruning. Additional loss curve derived by P2 Law are shown in Figure 131415 and 16 in Appendix G. As shown in these figures, the loss curve derived by P2 Law accurately aligns with all actual checkpoints, under all the three pruning methods, except for the width pruning on Llama-3.1-8B. As shown in Figure 2(c) and 8(c), for Llama-3.1-8B, we observe that depth pruning outperforms width pruning at similar pruning rates, which contrasts with the observations in other cases. This suggests that width pruning on Llama-3.1-8B may lead to anomalous performance, making our law unsuitable for this special scenario. We elaborate on this anomalous performance of width pruning on Llama-3.1-8B further in Appendix H.
@@ 4.3.1 Settings
Model size. We fit the P2 Law using checkpoints from smaller LLMs and validate it on checkpoints from larger LLMs, while maintaining the pruning rate during both fitting and prediction. Taking the Qwen-2.5 series models as an example, we fit the P2 Law using all checkpoints from Qwen-2.5-0.5B and Qwen-2.5-1.5B, and subsequently validate it with the actual checkpoints of Qwen-2.5-3B across three pruning rates. Due to the limited number of available actual loss curves for 2:4 semi-structured pruning, we did not conduct experiments for this pruning method.
@@
Pruning Rate. We fit the P2 Law using checkpoints from lower pruning rates and validate it using checkpoints from higher pruning rates, while keeping the model size constant during both fitting and prediction. Taking width pruning of the Qwen-2.5 series models as an example, we fit the P2 Law using checkpoints from these models at lower pruning rates (0.15 and 0.25) and then validate it with the actual checkpoints at a higher pruning rate of 0.35. Since there is no pruning rate in the 2:4 semi-structured pruning, we only explore the generalization ability on pruning rates under depth pruning and width pruning.
@@
Due to the anomaly of width pruning on Llama-3.1-8B (see Section 4.2), we exclude this model from generalization experiments.
4.3.2 Experimental Results
@@
Dataset Size Generalization. The evaluation results are shown in Table 3, and the loss curves of Qwen-2.5-3B (pruned by width pruning) derived by P2 Law are illustrated in Figure 7(a). Additional loss curves derived by P2 Law are provided in Figures 181920, and 21 in Appendix I. The results in Table 3 show that the loss curves derived by P2 Law accurately matches the validation checkpoints, indicating that the P2 Law generalizes well to larger dataset sizes.
@@
Model Size Generalization. The evaluation results are presented in Table 3, and the loss curves of Qwen-2.5-3B predicted by P2 Law (pruned by width pruning) are visualized in Figure 7(b). Additional loss curves predicted by P2 Law are shown in Figure 23 in Appendix I. As shown in Table 3, the P2 Law fitted on smaller LLMs performs poorly in $R^{2}$ and Huber loss when applied to larger models, indicating challenges in generalizing to larger, unseen models. However, the low ASD suggests it still captures the slope of the actual loss curve. This trend is also seen in Figure 23, where despite a gap between predicted and actual loss curves, the predicted and actual loss curves align in their downward trend after training stabilizes. This suggests P2 Law fitted from smaller LLMs can still predict the optimal computation cost point for larger LLMs, confirming its generalization feasibility.
@@
Pruning Rate Generalization. We present the generalization evaluations in Table 3 and illustrate the loss curves of Qwen-2.5 series models predicted by P2 Law (pruned by width pruning) in Figure 7(c). Additional loss curves predicted by P2 Law are provided in Figure 24 and 25 in Appendix I. As shown in the Figure 7(c) and Table 3, the values of different metrics indicate that the actual loss curves closely align with the predicted loss curves, suggesting that the P2 Law generalizes well to higher pruning rates.